OpenStudy (anonymous):

Solve: (x-2)/3 - (x-2)/5 = 2 I keep getting 13 when the answer is 7.

hartnn (hartnn):

can you show your steps ? i'll help you spot the error :)

OpenStudy (anonymous):

(5x+10)/15 - (3x-6)/15 = 2 (2x+4)/15 = 2 I multiplied both sides by 15 to get rid of the denominator 2x+4=30 2x=26 x=13

hartnn (hartnn):

shouldn't that be 5x-10 instead of 5x+10 ??? and actually , the correct answer is 17, not 7

hartnn (hartnn):

5(x-2) = 5x -10

hartnn (hartnn):

oh, its x+2 in the question...

hartnn (hartnn):

so, your mistake actually is (5x+10)/15 - (3x-6)/15 = (5x+10 - 3x -(-6)) /15

hartnn (hartnn):

5x+10 -3x +6 = 2x +16 you got 2x+4 instead

OpenStudy (anonymous):

I just got (5x+10 - 3x -(-6)) /15 incorrect. So it would be 2x+16 instead of 2x+4, which would end up giving me the 7 I need.